The rate at which a body cools also depends on its exposed surface area S. If S is a constant, then a modification of (2), given in Section 3.1, is dT/dt = kS(T-Tm), where
k < 0
and Tm is a constant. Suppose that two cups A and B are filled with coffee at the same time. Initially, the temperature of the coffee is 65° C. The exposed surface area of the coffee in cup B is twice the surface area of the coffee in cup A. After 30 min the temperature of the coffee in cup A is 38° C. If Tm = 21° C, then what is the temperature of the coffee in cup B after 30 min? (Round your answer to two decimal places.)

______ degrees Celsius
1
Expert's answer
2020-06-23T15:16:15-0400

"\frac {dT}{dt}" = kS(T-Tm) where k < 0

Here Tm= 21°C and S is exposed surface area.

Here "\frac {S_B}{S_A} = 2" , where SA and SB are the exposed surface area of A-cup and B-cup respectively.

Therefore

"\frac {dT}{(T - 21)}" = kSdt

Integrating,

"\int" "\frac {dT}{(T - 21)}" = "\int" kS dt

=> ln(T-21) = kSt + C, C is integration constant.

Temperature are being measured in °C and time in minute here.

When t = 0 , initial temperature T = 65

So, ln(65 - 21) = C

=> C = ln(44)

So ln(T-21) = kSt + ln(44 )

So ln(T-21) - ln(44) = kSt

=> ln "\frac {T-21}{44}" = kSt

For cup A, S = SA and T = 38 when t = 30

So equation for cup-A is

ln "\frac {38-21}{44}" = 30kSA

=> ln "\frac {17}{44}" = 30kSA •••••••••eq(1)

For cup-B, S = SB and T=? when t = 30

So equation for cup-B is

ln "\frac {T-21}{44}" = 30kSB

ln "\frac {T-21}{44}" = 60kSA as SB =2 SA ••••eq(2)

Comparing eq(2) and eq(1) we get

ln "\frac {T-21}{44}" = 2 ln "\frac {17}{44}"

=> ln "\frac {T-21}{44}" = ln ("\frac {17}{44}" )2

=> "\frac {T-21}{44}" = "(\frac {17}{44})^2"

=> T - 21 = "\frac {289}{44}"

=> T = 21 + 6.568

=> T = 27.568

The temperature of coffee in cup B is 27.57 °C ( correct up to two decimal places)